# ViCO: A Training Strategy towards Semantic Aware Dynamic High-Resolution

### v1

@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) (Zhu et al., 2025; Chen et al., 2024c; Xiaomi, 2025; Kwai Keye et al., 2025; Anthropic, 2025; DeepMind, 2025; Wang et al., 2024b) have demonstrated remarkable performance across a wide range of tasks, showing tremendous potential for real-world applications. Despite these advancements, MLLMs still suffer from the increased inference costs due to the additional vision tokens introduced by image inputs. Taking InternVL3.5 (Wang et al., 2025b) as an example, under its default configuration (Ye et al., 2023), each image is divided into up to 13 patches (including one thumbnail). Each patch is then represented by 256 visual tokens, resulting in a maximum of 3,328 tokens per image. In real-world scenarios such as document understanding or video comprehension, models are often required to process multiple images simultaneously. In such cases, the visual component will serve as the main body of the token sequence and constitutes the primary source of inference cost.
@@
The training procedure of ViCO consists of two stages: (1) Consistency Training: The model is trained to minimize the KL divergence between the responses conditioned on visual tokens with different compression rates. This encourages the model to generate accurate responses even when using highly compressed visual tokens, thereby improving its performance and robustness under compressed visual tokens. (2) Router Training: Although consistency training improves performance under high compression, inevitable information loss still causes a drop in performance. To mitigate this, we introduce ViR automatically selects the appropriate compression rate for each image patch. Those containing complex semantic information are represented using more tokens, while simpler patches are represented using fewer tokens. Compared with existing dynamic high-resolution strategies (Chen et al., 2024d; Wang et al., 2024a), which adjust the number of visual tokens based on image resolutions, our method further determines the number of visual tokens allocated for each image patch at the semantic level. This fine-grained control enables substantial efficiency gains with minimal performance loss.
To evaluate the effectiveness of our method, we conducted extensive experiments on benchmarks spanning OCR, document understanding, video understanding, and multi-image reasoning. Experimental results show that our method can reduce the number of visual tokens by up to 50% while preserving the model’s perception, reasoning, and OCR capabilities. As shown in Figure 1(b), our method maintains the original performance while improving the first token throughput of InternVL3.5 series (Wang et al., 2025b) across different model scales.
@@ 2 Related Work
With the development of large language models (Yang et al., 2025; OpenAI, 2025; Ouyang et al., 2022), multimodal large language models (MLLMs) have also made remarkable progress. To leverage LLMs and vision foundation models that have been pre-trained on unimodal datasets, a series of studies (Wang et al., 2024c; Liu et al., 2023a; d; Wang et al., 2025a; Li et al., 2023a) employ a connector to align the representational spaces of vision and language. This approach allows visual feature maps to be flattened and fed into LLMs as soft prompts, achieving strong performance through relatively low-cost incremental training. In addition, some works (Dubey et al., 2024; Luo et al., 2025) extend pre-trained LLMs by incorporating additional vision-language fusion layers. These layers enable the model to process tokens from different modalities with separate parameters, reducing the gradient conflicts across different modalities. However, the introduction of a large number of untrained parameters also brings additional training costs. More recently, several studies (Luo et al., 2024) have explored architectures without dedicated visual encoders. These models employ a unified Transformer to jointly process visual and textual information, eliminating the need for a separate vision encoder and fusion layer. Although MLLMs vary in their architectural designs, most of them adopt a dynamic high-resolution strategy (Chen et al., 2024c; d), which segments images into patches based on their resolution to enhance the model’s perceptual capabilities. However, this approach requires a large number of tokens to represent each image, thereby increasing the inference cost of MLLMs. In this paper, we propose a dynamic resolution strategy that is compatible with this paradigm. Our method introduces semantic-level adaptivity in determining the number of visual tokens needed to represent each image patch. This reduces the number of visual tokens and consequently the inference cost, while maintaining nearly the same performance.
@@
Improving the efficiency of large vision-language models (LVLMs) has drawn increasing attention, with visual token compression emerging as one of the most widely explored directions. Early approaches, such as LLaMA-VID (Li et al., 2024) and DeCo (Yao et al., 2024), aim to reduce redundancy in visual inputs through context tokens or adaptive pooling mechanisms, thereby lowering computational cost while retaining essential information. Similarly, MADTP (Cao et al., 2024) further enhances efficiency by identifying redundant tokens across different modalities and selectively removing them based on feature alignment, enabling more focused processing of relevant visual features. A number of lightweight, training-free methods have been proposed to reduce token redundancy without additional training. For instance, FastV (Chen et al., 2024a) prunes tokens in the LLM based on attention scores, removing low-attention tokens, while similar approaches such as SparseVLM (Zhang et al., 2024) and ToMe (Bolya et al., 2022) also reduce redundancy through token merging or selection without requiring retraining. While these strategies provide tangible computational benefits, their effectiveness can be limited on vision-sensitive tasks such as OCR, where fine-grained spatial and textual details are crucial. In such scenarios, overly aggressive token compression may lead to loss of essential visual information, highlighting the need for methods that carefully balance efficiency with the preservation of detailed visual representations.
@@ 3.1 Consistency Training
$\mathcal{L}_{\text{ViCO}}=\mathbb{E}_{\xi\sim\mathcal{R}}\Bigg[\frac{1}{N}\sum_{i=1}^{N}\mathrm{KL}\Big(\pi_{\theta_{ref}}\left(y_{i}\mid y_{<i},I\right)\;\Big\|\;\pi_{\theta_{policy}}\left(y_{i}\mid y_{<i},I_{\xi}\right)\Big)\Bigg],$ (1)
@@ 3.2 Router Training
$y_{i}^{\text{router}}=\begin{cases}0,&r_{i}<\tau\quad\text{(compression has negligible impact)}\\ 1,&r_{i}\geq\tau\quad\text{(compression has significant impact)},\end{cases}$ (3)
@@ 4.1 Experimental Settings
Benchmarks. We evaluate our approach on a diverse set of multimodal benchmarks. For OCR-related tasks, we use OCRBench (Liu et al., 2023c), ChartQA (Masry et al., 2022), and TextVQA (Singh et al., 2019), with InfoVQA (Mathew et al., 2022), DocVQA (Mathew et al., 2021), and AI2D (Kembhavi et al., 2016) additionally included in comparative and ablation studies. General multimodal tasks are evaluated on MMStar (Chen et al., 2024b), POPE (Li et al., 2023b), MME (Fu et al., 2023), MMBench V1.1 (Liu et al., 2023b), and RealWorldQA (Corp., 2024), while reasoning is assessed using the MMMU (Yue et al., 2023) benchmark. Multi-image understanding is measured on BLINK (Fu et al., 2024b), MMT-Bench (Ying et al., 2024), and MMIU (Ying et al., 2024), and video comprehension is tested on Video-MME (Fu et al., 2024a) under two settings: with subtitles and without subtitles.
@@ 4.5 Visualization of Image Routing
Results. As shown in Figure 3, our router is able to distinguish between semantically complex and simple patches. Regions containing salient objects, such as people, animals, or other key subjects, as well as areas with text carrying rich semantic information, are routed to lower compression rates and largely preserved at their original resolution. In contrast, relatively homogeneous background areas, which contain less critical information, are routed to higher compression rates. Importantly, being routed to higher compression does not imply that a patch is unimportant. Instead, it indicates that such information can be adequately represented with fewer resources. This selective compression strategy effectively reduces computational cost while allowing the model to focus more on patches that carry significant semantic content, ensuring that key visual information is retained for the LLM’s processing. Additional visualizations are provided in Appendix C.
@@ 4.6 Throughput Analysis
Table 4: Throughput of LLM under different token compression rates. Ori. refers to the original non-compressed model. We evaluate token compression rates of 25%, 50%, and 75% to measure the speedup. All values are reported as speedup factors ($\times$) relative to the non-compressed baseline.
@@
Settings. We deploy our ViCO models using the LMDeploy (Contributors, 2023) framework and evaluate different model scales, namely 8B, 30B, 38B, and 241B, under varying image–text ratios and compression rates. Using a real deployment framework ensures that the evaluation more accurately reflects practical inference performance. For throughput simulation, each request contains 10 visual patches with text inputs of length 512 or 64 tokens. The LLM’s first-token throughput is measured over 2000 requests, which are sent concurrently to the API using 32 threads, and the results are reported as relative speedup compared to the non-compressed baseline. We set the tensor parallelism of the 241B model to 8, and to 1 for all other models.
@@ 5 Conclusion
In this work, we propose Visual Consistency Learning (ViCO), which enables the model to represent images of varying semantic complexities using different numbers of vision tokens. Compared with existing dynamic high-resolution strategies, which adjust the number of visual tokens based on image resolutions, our method dynamically adapts the number of visual tokens according to semantic complexity. Experimental results demonstrate that our method can reduce the number of vision tokens by up to 50% while maintaining the model’s perception, reasoning, and OCR capabilities. Additional ablation studies further validate the effectiveness of each proposed module. By making decisions based on visual semantics, ViCO enables the model to efficiently focus computation on the most informative regions of an image, providing insights for future research on adaptive visual representations. We hope this work will contribute to the development of more efficient MLLMs. The code and models will be released to facilitate future research.
Acknowledgments
This work is supported by the National Key R&D Program of China (NO. 2022ZD0161302).
ETHICS STATEMENT
This study complies with the ICLR ethical guidelines and adheres to the principles of responsible research. We confirm that no personally identifiable, sensitive, or potentially harmful data were utilized. We have considered the potential impact of our methods and believe that they advance scientific understanding without causing any foreseeable harm.
REPRODUCIBILITY STATEMENT
We will open-source our code and model weights on platforms such as GitHub and Hugging Face, including methodological details and experimental settings, to ensure the reproducibility of our methods.
Appendix A THE USE OF LARGE LANGUAGE MODELS (LLMS)
Large language models (LLMs) were only employed for light linguistic refinement, such as polishing sentences and correcting grammatical errors. They were not involved in the formulation or development of the core ideas presented in this article.
Appendix B Performance of ViCO on small-scale models
@@
Table 5: Performance of ViCO on InternVL3.5 small-scale models across general benchmarks. When calculating the overall score, MME is normalized from 0–2800 to 0–100. RWQA refers to RealWorldQA.
Model OCRBench ChartQA TextVQA MMStar POPE MME MMBench V1.1 RWQA Avg.
@@
Ratio of Compressed Patches 68% 36% 60% 61% 77% 73% 71% 66% 64.0%
Table 6: Performance of ViCO on InternVL3.5 small-scale models across multi-image and video benchmarks. For Video-MME, both the subtitle and non-subtitle settings are evaluated using 64-frame inputs.
Model BLINK MMT-Bench MMIU Video-MME Avg.
@@
Ratio of Compressed Patches 69% 58% 48% 73% 73% 64.2%

### v2

@@ 1 Introduction
Recent advancements in Multimodal Large Language Models (MLLMs) [44, 6, 37, 16, 1, 10, 33] have demonstrated remarkable performance across a wide range of tasks, showing tremendous potential for real-world applications. Despite these advancements, MLLMs still suffer from the increased inference costs due to the additional vision tokens introduced by image inputs. Taking InternVL3.5 [36] as an example, under its default configuration [40], each image is divided into up to 13 patches (including one thumbnail). Each patch is then represented by 256 visual tokens, resulting in a maximum of 3,328 tokens per image. In real-world scenarios such as document understanding or video comprehension, models are often required to process multiple images simultaneously. In such cases, the visual component will serve as the main body of the token sequence and constitutes the primary source of inference cost.
@@
The training procedure of ViCO consists of two stages: (1) Consistency Training: The model is trained to minimize the KL divergence between the responses conditioned on visual tokens with different compression rates. This encourages the model to generate accurate responses even when using highly compressed visual tokens, thereby improving its performance and robustness under compressed visual tokens. (2) Router Training: Although consistency training improves performance under high compression, inevitable information loss still causes a drop in performance. To mitigate this, we introduce ViR, which automatically selects the appropriate compression rate for each image patch. Those containing complex semantic information are represented using more tokens, while simpler patches are represented using fewer tokens. Compared with existing dynamic high-resolution strategies [7, 32], which adjust the number of visual tokens based on image resolutions, our method further determines the number of visual tokens allocated for each image patch at the semantic level. This fine-grained control enables substantial efficiency gains with minimal performance loss.
To evaluate the effectiveness of our method, we conducted extensive experiments on benchmarks spanning OCR, document understanding, video understanding, and multi-image reasoning. Experimental results show that our method can reduce the number of visual tokens by up to 50% while preserving the model’s perception, reasoning, and OCR capabilities. As shown in Figure 1(b), our method maintains the original performance while improving the first token throughput of InternVL3.5 series [36] across different model scales.
@@ 2 Related Work
With the development of large language models [38, 29, 30], multimodal large language models (MLLMs) have also made remarkable progress. To leverage LLMs and vision foundation models that have been pre-trained on unimodal datasets, a series of studies [34, 20, 23, 35, 17] employ a connector to align the representational spaces of vision and language. This approach allows visual feature maps to be flattened and fed into LLMs as soft prompts, achieving strong performance through relatively low-cost incremental training. In addition, some works [11, 25] extend pre-trained LLMs by incorporating additional vision-language fusion layers. These layers enable the model to process tokens from different modalities with separate parameters, reducing the gradient conflicts across different modalities. However, the introduction of a large number of untrained parameters also brings additional training costs. More recently, several studies [24] have explored architectures without dedicated visual encoders. These models employ a unified Transformer to jointly process visual and textual information, eliminating the need for a separate vision encoder and fusion layer. Although MLLMs vary in their architectural designs, most of them adopt a dynamic high-resolution strategy [6, 7], which segments images into patches based on their resolution to enhance the model’s perceptual capabilities. However, this approach requires a large number of tokens to represent each image, thereby increasing the inference cost of MLLMs. In this paper, we propose a dynamic resolution strategy that is compatible with this paradigm. Our method introduces semantic-level adaptivity in determining the number of visual tokens needed to represent each image patch. This reduces the number of visual tokens and consequently the inference cost, while maintaining nearly the same performance.
@@
Improving the efficiency of large vision-language models (LVLMs) has drawn increasing attention, with visual token compression emerging as one of the most widely explored directions. Early approaches, such as LLaMA-VID [19] and DeCo [39], aim to reduce redundancy in visual inputs through context tokens or adaptive pooling mechanisms, thereby lowering computational cost while retaining essential information. Similarly, MADTP [3] further enhances efficiency by identifying redundant tokens across different modalities and selectively removing them based on feature alignment, enabling more focused processing of relevant visual features. A number of lightweight, training-free methods have been proposed to reduce token redundancy without additional training. For instance, FastV [5] prunes tokens in the LLM based on attention scores, removing low-attention tokens, while similar approaches such as SparseVLM [43] and ToMe [2] also reduce redundancy through token merging or selection without requiring retraining. While these strategies provide tangible computational benefits, their effectiveness can be limited on vision-sensitive tasks such as OCR, where fine-grained spatial and textual details are crucial. In such scenarios, overly aggressive token compression may lead to loss of essential visual information, highlighting the need for methods that carefully balance efficiency with the preservation of detailed visual representations.
@@ 3.1 Consistency Training
$\mathcal{L}_{\text{ViCO}}$ $=\mathbb{E}_{\xi\sim\mathcal{R}}\Bigg[\frac{1}{N}\sum_{i=1}^{N}\mathrm{KL}\Big(\pi_{\theta_{ref}}\!\left(y_{i}\mid y_{<i},I\right)$ (1)
$\qquad\qquad\Big\|\;\pi_{\theta_{policy}}\!\left(y_{i}\mid y_{<i},I_{\xi}\right)\Big)\Bigg],$
@@ 3.2 Router Training
$y_{i}^{\text{router}}=\begin{cases}0,&r_{i}<\tau\quad\text{(low compression impact)}\\ 1,&r_{i}\geq\tau\quad\text{(high compression impact)}\end{cases},$ (3)
@@ 4.1 Experimental Settings
Benchmarks. We evaluate our approach on a diverse set of multimodal benchmarks. For OCR-related tasks, we use OCRBench [22], ChartQA [26], and TextVQA [31], with InfoVQA [28], DocVQA [27], and AI2D [15] additionally included in comparative and ablation studies. General multimodal tasks are evaluated on MMStar [4], POPE [18], MME [12], MMBench V1.1 [21], and RealWorldQA [9], while reasoning is assessed using the MMMU [42] benchmark. Multi-image understanding is measured on BLINK [14], MMT-Bench [41], and MMIU [41], and video comprehension is tested on Video-MME [13] under two settings: with subtitles and without subtitles.
@@ 4.5 Visualization of Image Routing
Results. As shown in Figure 3, our router is able to distinguish between semantically complex and simple patches. Regions containing salient objects, such as people, animals, or other key subjects, as well as areas with text carrying rich semantic information, are routed to lower compression rates and largely preserved at their original resolution. In contrast, relatively homogeneous background areas, which contain less critical information, are routed to higher compression rates. Importantly, being routed to higher compression does not imply that a patch is unimportant. Instead, it indicates that such information can be adequately represented with fewer resources. This selective compression strategy effectively reduces computational cost while allowing the model to focus more on patches that carry significant semantic content, ensuring that key visual information is retained for the LLM’s processing.
@@ 4.6 Throughput Analysis
Table 4: Throughput of LLM under different token compression rates. Ori. refers to the original non-compressed model. We evaluate token compression rates of 25%, 50%, and 75% to measure the resulting speedup. All values are reported as speedup factors ($\times$) relative to the non-compressed baseline.
@@
Settings. We deploy our ViCO models using the LMDeploy [8] framework and evaluate different model scales, namely 8B, 30B, 38B, and 241B, under varying image–text ratios and compression rates. Using a real deployment framework ensures that the evaluation more accurately reflects practical inference performance. For throughput simulation, each request contains 10 visual patches with text inputs of length 512 or 64 tokens. The LLM’s first-token throughput is measured over 2000 requests, which are sent concurrently to the API using 32 threads, and the results are reported as relative speedup compared to the non-compressed baseline. We set the tensor parallelism of the 241B model to 8, and to 1 for all other models.
@@ 5 Conclusion
In this work, we propose Visual Consistency Learning (ViCO), which enables the model to represent images of varying semantic complexities using different numbers of vision tokens. Compared with existing dynamic high-resolution strategies, which adjust the number of visual tokens based on image resolutions, our method dynamically adapts the number of visual tokens according to semantic complexity. Experimental results demonstrate that our method can reduce the number of vision tokens by up to 50% while maintaining the model’s perception, reasoning, and OCR capabilities. Additional ablation studies further validate the effectiveness of each proposed module. By making decisions based on visual semantics, ViCO enables the model to efficiently focus computation on the most informative regions of an image, providing insights for future research on adaptive visual representations. We hope this work will contribute to the development of more efficient MLLMs by reducing inference-time computational cost. The code and models will be released to facilitate future research.
Supplementary Material
Appendix A Performance of ViCO on small-scale models
@@
Table 5: Performance of ViCO on InternVL3.5 small-scale models across general benchmarks. MME is normalized from 0–2800 to 0–100. Abbreviations: OCR = OCRBench, Chart = ChartQA, Text = TextVQA, MMB = MMBench, RWQA = RealWorldQA, Ratio = Ratio of Compressed Patches.
Model OCR Chart Text MMStar POPE MME MMB V1.1 RWQA Avg.
@@
[…]
Table 6: Performance of ViCO on InternVL3.5 small-scale models across multi-image and video benchmarks. For Video-MME, both the subtitle and non-subtitle settings are evaluated using 64-frame inputs. Abbreviations: MMT = MMT-Bench, Ratio = Ratio of Compressed Patches.
Model BLINK MMT MMIU Video-MME Avg.
@@
[…]
Appendix B Discussion on Additional Inference Latency
In this work, we introduce a new Router structure. Specifically, for models with 1B, 2B, 4B, 8B, 14B, and 30B parameters, the Router has approximately 50M parameters, while for 38B and 241B models, it has around 490M parameters, which is very small relative to the overall model size. Additionally, the Attention Pooling module scales linearly with sequence length in terms of computational complexity. Therefore, the theoretical computational overhead introduced by these additional structures is minimal. Moreover, as demonstrated in our experiments in the main text, these structures achieve notable acceleration in practical deployment scenarios. In summary, the added components do not introduce significant extra computational cost or inference latency.